# Boosting Adversarial Transferability with Spatial Adversarial Alignment

### v1

@@ 1 Introduction
Deep neural networks (DNNs) have been successfully and extensively deployed across security-sensitive applications, including autonomous driving [46, 51, 50], facial verification [53, 52, 34], and video surveillance [17, 54, 18, 14]. However, DNNs exhibit considerable vulnerability to adversarial examples [13, 27, 12, 3, 4, 6, 5], where imperceptible perturbations are introduced into natural images, leading models to produce incorrect predictions. In real-world applications, DNNs are typically concealed from user access, necessitating adversaries to generate adversarial examples within a black-box setting, where no knowledge of the target model’s parameters or architecture is available. Adversarial transferability plays a crucial role in black-box settings as it allows adversaries to effectively compromise target models by employing adversarial examples generated on surrogate models. In black-box settings, adversarial transferability plays a crucial role, which enables adversaries to leverage adversarial examples crafted on surrogate models to effectively attack target models. Thus, generating highly transferable adversarial examples is instrumental in uncovering and understanding the vulnerabilities within DNNs, drawing substantial attention in recent research.
@@
Cross-model transferability has been extensively studied for CNNs [9, 45, 10]. Highly transferable adversarial examples are usually based on advanced optimization [9, 22, 37] and data augmentation [45, 10, 25]. The principle is to alleviate the overfitting of adversarial examples on surrogate models, determining whether the attack can be successfully transferred to the target models. In addition, some model modification methods [44, 15], such as amplifying the gradient on skip connections (the structure in ResNet [16]), can also improve transferability. However, few works explore adversarial transferability on Vision Transformer (ViT) [11] and the performance of existing work extending CNN to ViT is poor due to significant structural differences. Specifically, ViT flattens the image into a sequence of patch tokens and employs multi-head self-attention to capture global relationships among the patches. In contrast, CNNs typically consist of stacked convolutional layers that learn feature relationships progressively through downsampling. Therefore, [43] first empirically analyzes the structure of ViT and propose PNA and PatchOut [43], but there is still much room for improvement in cross-architecture transferability.
@@
In this paper, we argue that unique structural features are critical to cross-architecture adversarial transferability. Given a dataset, various models tend to exhibit analogous decision boundaries [23], arising from their ability to learn similar features. If we can obtain a surrogate model whose features are similar to those of models with different architectures, then the resulting adversarial perturbation can be transferable across different models. A recent technique known as Model Alignment (MA) [26] employs an alignment loss to minimize prediction divergences between surrogate models and witness models, thereby indirectly facilitating the extraction of features that are similarly represented by the witness model. However, directly applying MA to black-box attacks may lead to the degradation of cross-architecture transferability. The main reasons are: (i) Features are not aligned in space. MA only uses the final prediction of the model, but in fact, the spatial features of ViT and CNN are different. It is difficult to directly constrain the similarity of features only by the final logits. (ii) Features are not aligned from the perspective of adversarial features. In addition to the features of clean images, the features of adversarial examples also have similarities across different models and need to be considered.
@@
To overcome these challenges and enhance transferability, we propose a technique called Spatial Adversarial Alignment (SAA), which utilizes an alignment loss from the perspective of spatial and adversarial features and incorporates a witness model to refine the surrogate model. SAA consists of two key parts: spatial-aware alignment and adversarial-aware alignment. In the spatial-aware alignment, in addition to aligning on the final global features, we also focus on the features of local regions. We make local features of CNNs by position to align ViTs’ embeddings at the same position. In the adversarial-aware alignment, we introduce a self-adversarial strategy, which constructs adversarial examples so that the model can learn the differences between different architectures in adversarial features, thereby enabling the model to further capture more common features. Aligned surrogate models by SAA provide promising adversarial transferability and can be seamlessly integrated with existing transfer attacks. Our contributions can be summarized as follows:
  • We reveal for the first time the importance of spatial and adversarial features for cross-architecture transferability.
@@
  • We propose Spatial Adversarial Alignment (SAA), which leverages a witness model to fine-tune the surrogate model via spatial-aware and adversarial-aware alignment to generate highly transferable adversarial examples.
@@
  • Experiments on 6 CNNs and 4 ViTs show that SAA has state-of-the-art adversarial transferability, especially in cross-model transferability. Compared with MA, on ResNet50, the transferability from CNN to ViT is improved by 25.5-39.1%.
2 Related Work
2.1 Transfer attacks on CNNs
2.2 Transfer attacks on ViTs
@@ 3.1 Preliminaries
In this paper, we focus on the image classification task on DNNs. Let $f_{\theta}(\cdot)$ represent a DNN-based classifier with parameters $\theta$. We denote the clean image as $x$ and its corresponding ground-truth label as $y$. Following [9, 45, 10], we evaluate the adversarial transferability under untargeted adversarial attacks with $l_{\infty}$ norm. Therefore, the goal of transfer attacks is to add an adversarial perturbation to the clean image $x$ based on the information of the surrogate model $f_{\theta_{s}}(\cdot)$ to obtain the adversarial example $x_{adv}$, so that $f_{\theta_{s}}(x_{adv})\neq y$ subject to the constraint that $||x_{adv}-x||_{\infty}\leq~{}\epsilon$. In the black-box setting, no information about the target model—such as its architecture, weights, or gradients—is accessible. Therefore, adversarial examples are generated solely by utilizing a surrogate model $f_{\theta_{s}}(\cdot)$, leveraging their transferability to deceive the target model $f_{\theta_{t}}(\cdot)$.
@@ 3.2 Spatial Adversarial Alignment (SAA)
Spatial-aware Alignment. The purpose of spatial-aware alignment is to make the surrogate and witness models more consistent in the feature space. Naturally, the most intuitive approach to aligning the feature distributions of two models is to minimize the distance between their final outputs. However, when the models exhibit significant architectural differences, ensuring output similarity alone is insufficient to achieve alignment in intermediate features. In black-box attacks, where the details of the target model’s architecture are unknown, this issue becomes more pronounced. Taking the challenging scenario of CNN to ViT as an example, their intermediate layer features differ substantially in semantic levels. This discrepancy arises primarily from differences in receptive fields, stacking methodologies, and normalization techniques between CNNs and ViTs. Therefore, relying solely on output alignment for model fine-tuning indirectly captures some common features, but this approach can, in certain cases, result in degraded transferability, as observed in methods like Model Alignment (MA) [26].
Therefore, in addition to aligning on the final global features, we also need focus on the features of local regions. For ease of understanding, we define the global features $f_{\theta}(x)$ as the logits of the model corresponding to the input $x$. For CNNs, it is the output of features by the last layer. For ViTs, it refers to the final embedding of the [CLS] token after the MLP block. First, we perform alignment at the global feature level by defining an alignment loss between the surrogate model and witness model at the output layer:
@@
Next, we align the models at the local feature level. Here, we define $z_{\theta}(x)$ as local features. For CNNs, $z_{\theta}(x)^{B\times C\times H\times W}$ is the features of the last layer before passing through the global average pooling layer. For ViTs, $z_{\theta}(x)^{B\times C^{\prime}\times H\times W}$***Generally, ViT’s patch embeddings $z(x)$ is $(B,N,C’)$ ’ ) by default. We first transform it to $(B,C^{\prime},H’,W’)$ ’ ), where $N=H^{\prime}\times W^{\prime}$. Then, we perform an adaptive pooling operation to transform it to $(B,C^{\prime},H,W)$. is the embeddings of the patch tokens after passing through the MLP except for the [CLS] token, where each patch token corresponds to a specific spatial region in the input. Each spatial position $(h,w)$ within this feature map is treated as a distinct local region and the feature for each local region is $z_{\theta}^{[q]}(x)$, where $q=\{1,2,...,H\times W\}$. Let $z_{\theta_{s}}^{[q]}(x)$ and $z_{\theta_{w}}^{[q]}(x)$ denote the local prediction (logits) associated with each local region $q$ for the surrogate model and witness model, respectively. Next, we compute the pseudo-labels of the local region $q$ based on the local features of the witness models. We denote this pseudo-label as $\hat{y}^{[q]}_{\theta_{w}}$, which is obtained by taking the $\arg\max$ over the logits from the witness model: $\hat{y}^{[q]}_{\theta_{w}}=\arg\max(f_{\theta_{w}}(z_{\theta_{w}}^{[q]}(x)))$. Then, we use this pseudo-label to supervise the learning of the local feature of the surrogate model. To achieve local alignment, we minimize the divergence between the logits of corresponding local regions, so the local alignment loss is expressed as:
$\small\mathcal{L}_{local}(x;\theta_{s})=\frac{1}{HW}\sum^{HW}_{q=1}D_{\mathrm{CE}}(z_{\theta_{s}}^{[q]}(x),\arg\max(f_{\theta_{w}}(z_{\theta_{w}}^{[q]}(x)))),$ (2)
@@
Adversarial-aware Alignment. The relationship between features and adversarial vulnerability is highly significant. Some hypotheses [47, 30] propose that adversarial examples possess distinct feature distributions compared to normal examples, which may inherently predispose models to adversarial vulnerability—a notion supported by several studies [1, 25]. Beyond normal examples, learning adversarial features may offer a way to capture shared features between surrogate models and witness models. Furthermore, [10] suggests that models trained with adversarial examples focus on more discriminative regions within images, displaying feature recognition patterns distinct from those of normally trained models. Thus, adversarial examples play a crucial role in achieving model alignment.
@@
Spatial-adversarial alignment facilitates the alignment of the surrogate model with a collection of witness models to further improve the adversarial transferability. This process involves utilizing a set of witness models, denoted as $\Theta$, where the cardinality of the set is represented by $|\Theta|$. Therefore, the parameter update rule for the surrogate model, based on stochastic gradient descent (SGD), can be expressed as follows:
$\theta_{s}^{(t+1)}=\theta_{s}^{(t)}-\eta\cdot\frac{1}{|\mathcal{B}||\Theta|}\sum\limits_{x\in\mathcal{B}}\sum\limits_{\theta_{w}\in\Theta}\nabla_{\theta_{s}^{(t)}}\mathcal{L}_{SAA}(x;\theta_{s}),$ (7)
where $t$ is the epoch, $\eta$ is the learning rate and $\mathcal{B}$ means the mini-batch samples.
@@ 3.3 A Close Look at SAA
To verify whether SAA significantly improves the spatial and adversarial features after model alignment, we conduct quantitative and qualitative analyses based on the models before and after alignment. We randomly sample 100 images from ImageNet val and then compute the cosine similarity between the global features of the surrogate models before and after applying SAA with different witness models. Table 1 demonstrates that, whether for clean images or adversarial examples (generated by SSA-DI-TI-MI), the feature similarity improves after alignment. Notably, when the surrogate model is ViT-B, the improvement in similarity is even more pronounced. This result suggests that, after applying SAA, the aligned surrogate models effectively capture features shared with the witness model, providing strong evidence of the alignment’s success.
@@
Then, we use Grad-CAM [32]’s heatmaps to simulate the feature distribution of the model, as shown as Figure 2. For the clean inputs (first four cols), the heatmaps generated by the unaligned surrogate model (2-nd col) primarily focus on local regions of the object. In contrast, the aligned surrogate model (4-th col) heatmaps demonstrate more diffuse attention spread across the entire object, similar to that of the witness model (ViT-B, 3-rd col), which shows aligned surrogate models learn the common spatial features. For adversarial examples (last four cols), the 5-th and 7th cols display the heatmaps of adversarial examples generated by the unaligned and aligned surrogate models, respectively. The 6-th and 8-th show the witness model’s responses to these adversarial examples. Notably, the adversarial examples generated by the unaligned surrogate model fail to effectively transfer to the witness model (6-th col) due to still focusing on the target subject, indicating limited cross-model transferability. In contrast, adversarial examples generated by the aligned surrogate model (7-th col) successfully transferred to the witness model (8-th col) as the features are spread out, demonstrating enhanced cross-model transferability achieved through SAA.
@@ 4.1 Experimental Setup
Datasets. Our experiments utilize the ImageNet-compatible dataset [21], a widely adopted subset containing 1,000 images from the ImageNet validation set [7]. This dataset is commonly used in adversarial robustness studies, such as those in [45, 9, 10].
Models. To assess the adversarial transferability of different network architectures, we focus on convolutional neural networks (CNNs) and vision transformers (ViTs) as the target models. For CNNs, we select the typically trained ResNet-18 (Res18), ResNet-50 (Res50) and ResNet-101 (Res101) [16], VGG-19 [33], DenseNet-121 (DN121) [19], and Inception-v3 (Inc-v3) [35]. For ViTs, we evaluate the Vision Transformer (ViT-B) [11], Swin Transformer (Swin-B) [24], Pyramid Vision Transformer (PVT-v2) [38], and MobileViT-s (MobViT) [28].
Metric. Adversarial transferability is quantified by calculating the average attack success rate (Avg. ASR, %) across target models (excluding the surrogate model), with a higher success rate signifying enhanced transferability. In the paper, ’n/a’ is defined as the average attack success rate obtained by generating adversarial examples using the surrogate model without any alignment, serving as a baseline.
@@
Surrogate Witness Attack Target Model Avg. ASR (%)
@@
Implementation Details. In our experiments, we select the MI [9] attack as the baseline for generating adversarial examples with high transferability, as it is widely recognized within the field of adversarial transferability [10, 45, 39, 25, 40, 37, 22, 42, 44, 43]. For MI, we set the perturbation magnitude $\epsilon=16$, perform 10 iterations, with a step size of $\frac{16}{10}=1.6$, and use a momentum $\mu=1$. During the Spatial Adversarial Alignment, all surrogate models are fine-tuned for one epoch using stochastic gradient descent (SGD) with a momentum of 0.9, and no learning rate adjustments are applied. It is important to note that no additional data is used for fine-tuning, as it relies solely on the same training data used for both the surrogate and witness models. The settings for Model Alignment (MA) [26] are consistent with the parameters specified in the original paper.
Table 3: SAA has stronger adversarial transferability after being compatible with existing transfer attacks.

### v2

@@ 1 Introduction
Deep neural networks (DNNs) have been successfully and extensively deployed across security-sensitive applications liu2024a , including autonomous driving yang2023aide ; Zhang_2024_CVPR ; zhang2024efficient , facial verification 10448479 ; ZhengLWWMLDW24 ; sun2024rethinking ; ci2024ringid , and video surveillance hong2023lvos ; zhou23nips ; hong2024onetracker ; guo2024x . However, DNNs exhibit considerable vulnerability to adversarial examples fgsm ; pgd ; autoattack ; chen2022shape ; chen2022towards ; chen2023aca ; chen2023devopatch ; liu2022imperceptible ; liu2023point ; liu2022efficient , where imperceptible perturbations are introduced into natural images, leading models to produce incorrect predictions. In real-world applications, DNNs are typically concealed from user access, necessitating adversaries to generate adversarial examples within a black-box setting, where no knowledge of the target model’s parameters or architecture is available. Adversarial transferability mi ; jiang2023exploring ; WANG2024124757 plays a crucial role in black-box settings as it allows adversaries to effectively compromise target models by employing adversarial examples generated on surrogate models. In black-box settings, adversarial transferability plays a crucial role, which enables adversaries to leverage adversarial examples crafted on surrogate models to effectively attack target models. Thus, generating highly transferable adversarial examples is instrumental in uncovering and understanding the vulnerabilities within DNNs, drawing substantial attention in recent research.
@@
Cross-model transferability has been extensively studied for CNNs mi ; xie2019improving ; dong2019evading . Highly transferable adversarial examples are usually based on advanced optimization mi ; sini ; WANG2024124757 and data augmentation xie2019improving ; dong2019evading ; ssa2022 . The principle is to alleviate the overfitting of adversarial examples on surrogate models, determining whether the attack can be successfully transferred to the target models. In addition, some model modification methods sgm ; linbp2020 , such as amplifying the gradient on skip connections (the structure in ResNet resnet ), can also improve transferability. However, few works explore adversarial transferability on Vision Transformer (ViT) vit and the performance of existing work extending CNN to ViT is poor due to significant structural differences. Specifically, ViT flattens the image into a sequence of patch tokens and employs multi-head self-attention to capture global relationships among the patches. In contrast, CNNs typically consist of stacked convolutional layers that learn feature relationships progressively through downsampling. Therefore, pna first empirically analyzes the structure of ViT and propose PNA and PatchOut pna , but there is still much room for improvement in cross-architecture transferability.
@@
In this paper, we argue that unique structural features are critical to cross-architecture adversarial transferability. Given a dataset, various models tend to exhibit analogous decision boundaries decisions , arising from their ability to learn similar features. If we can obtain a surrogate model whose features are similar to those of models with different architectures, then the resulting adversarial perturbation can be transferable across different models. A recent technique known as Model Alignment (MA) ma2025improving employs an alignment loss to minimize prediction divergences between surrogate models and witness models, thereby indirectly facilitating the extraction of features that are similarly represented by the witness model. However, directly applying MA to black-box attacks may lead to the degradation of cross-architecture transferability. The main reasons are: (i) Features are not aligned in space pna . MA only uses the final prediction of the model, but in fact, the spatial features of ViT and CNN are different zhou2022understanding ; Wang_2024_CVPR . It is difficult to directly constrain the similarity of features only by the final logits. (ii) Features are not aligned from the perspective of adversarial features. In addition to the features of clean images, the features of adversarial examples also have similarities across different models and need to be considered.
@@
To overcome these challenges and enhance transferability, we propose a technique called Spatial Adversarial Alignment (SAA), which utilizes an alignment loss from the perspective of spatial and adversarial features and incorporates a witness model to refine the surrogate model. SAA consists of two key parts: spatial-aware alignment and adversarial-aware alignment. In the spatial-aware alignment, in addition to aligning on the final global features, we also focus on the features of local regions. We make local features of CNNs by position to align ViTs’ embeddings at the same position. In the adversarial-aware alignment, we introduce a self-adversarial strategy, which constructs adversarial examples so that the model can learn the differences between different architectures in adversarial features, thereby enabling the model to further capture more common features. Aligned surrogate models by SAA provide promising adversarial transferability and can be seamlessly integrated with existing transfer attacks. In addition, we further summarize the empirical guidance for the selection of surrogate and witness models in SAA. Our contributions can be summarized as follows:
[…]
@@
$\bullet$ We propose Spatial Adversarial Alignment (SAA), which leverages a witness model to fine-tune the surrogate model via spatial-aware and adversarial-aware alignment to generate highly transferable adversarial examples. In addition, we further summarize the empirical guidance for the model selection in SAA.
@@
$\bullet$ Experiments on 6 CNNs and 4 ViTs show that SAA has state-of-the-art adversarial transferability, especially in cross-model transferability. Compared with MA, on ResNet50, the transferability from CNN to ViT is improved by 25.5-39.1%.
@@ 2.1 Preliminaries
In this paper, we focus on the image classification task on DNNs. Let $f_{\theta}(\cdot)$ represent a DNN-based classifier with different network parameters $\theta$. We denote the clean image as $x$ and its corresponding ground-truth label as $y$. Following mi ; xie2019improving ; dong2019evading , we evaluate the adversarial transferability under untargeted adversarial attacks with $l_{\infty}$ norm. Therefore, the goal of transfer attacks is to add an adversarial perturbation to the clean image $x$ based on the information of the surrogate model $f_{\theta_{s}}(\cdot)$ to obtain the adversarial example $x_{adv}$ fgsm , so that $f_{\theta_{s}}(x_{adv})\neq y$ subject to the constraint that $x_{adv}-x||_{\infty}\leq~\epsilon$.
In the black-box setting, no information about the target model—such as its architecture, weights, or gradients—is accessible. Therefore, adversarial examples are generated solely by utilizing a surrogate model $f_{\theta_{s}}(\cdot)$, leveraging their transferability to deceive the target model $f_{\theta_{t}}(\cdot)$.
@@ 2.2 Spatial Adversarial Alignment (SAA)
Spatial-aware Alignment. The purpose of spatial-aware alignment is to make the surrogate and witness models more consistent in the feature space. Naturally, the most intuitive approach to aligning the feature distributions of two models is to minimize the distance between their final outputs ma2025improving . However, when the models exhibit significant architectural differences, ensuring output similarity alone is insufficient to achieve alignment in intermediate features. In black-box attacks, where the details of the target model’s architecture are unknown, this issue becomes more pronounced. Taking the challenging scenario of CNN to ViT as an example, their intermediate layer features differ substantially in semantic levels zhou2022understanding . This discrepancy arises primarily from differences in receptive fields, stacking methodologies, and normalization techniques between CNNs and ViTs. Therefore, relying solely on output alignment for model fine-tuning indirectly captures some common features, but this approach can, in certain cases, result in degraded transferability, as observed in methods like Model Alignment (MA) ma2025improving .
Therefore, in addition to aligning on the final global features, we also need to focus on the features of local regions. For ease of understanding, we define the global features $f_{\theta}(x)$ as the logits of the model corresponding to the input $x$. For CNNs, it is the output of features by the last layer. For ViTs, it refers to the final embedding of the [CLS] token after the MLP block. First, we perform alignment at the global feature level by defining an alignment loss between the surrogate model and witness model at the output layer:
@@
[…]
@@
Adversarial-aware Alignment. The relationship between features and adversarial vulnerability is highly significant. Some hypotheses c-gsp ; nie2022DiffPure propose that adversarial examples possess distinct feature distributions compared to normal examples, which may inherently predispose models to adversarial vulnerability—a notion supported by several studies cfm ; ssa2022 . Beyond normal examples, learning adversarial features may offer a way to capture shared features between surrogate models and witness models. Furthermore, dong2019evading suggests that models trained with adversarial examples focus on more discriminative regions within images, displaying feature recognition patterns distinct from those of normally trained models. Thus, adversarial examples play a crucial role in achieving model alignment.
@@
Spatial-adversarial alignment facilitates the alignment of the surrogate model with the witness model to further improve the adversarial transferability. The parameter update rule for the surrogate model, based on stochastic gradient descent (SGD), can be expressed as follows:
$\theta_{s}^{(t+1)}=\theta_{s}^{(t)}-\eta\cdot\frac{1}{|\mathcal{B}|}\sum\limits_{x\in\mathcal{B}}\nabla_{\theta_{s}^{(t)}}\mathcal{L}_{SAA}(x;\theta_{s}),$ (7)
where $t$ is the epoch, $\eta$ is the learning rate and $\mathcal{B}$ means the mini-batch samples. Please refer to Algorithm 1 for the detailed loss calculation process.
@@ 2.3 A Close Look at SAA
To verify whether SAA significantly improves the spatial and adversarial features after model alignment, we conduct quantitative and qualitative analyses based on the models before and after alignment. We randomly sample 100 images from ImageNet val and then compute the cosine similarity between the global features of the surrogate models before and after applying SAA with different witness models. Table 1 shows that, whether for clean images or adversarial examples (generated by SSA-DI-TI-MI), the feature similarity improves after alignment. Notably, when the surrogate model is ViT-B, the improvement in similarity is even more pronounced. In addition, the feature gap between Res50 and Swin-B is indeed significant, but even so, SAA achieves a $\frac{0.0551-0.0369}{0.0369}=49.3\%$ improvement in similarity. This result suggests that, after applying SAA, the aligned surrogate models effectively capture features shared with the witness model, providing strong evidence of the alignment’s success.
@@
Then, we use Grad-CAM gradcam ’s heatmaps to simulate the feature distribution of the model, as shown as Figure 2. For the clean inputs (first four cols), the heatmaps generated by the unaligned surrogate model (2-nd col) primarily focus on local regions of the object. In contrast, the aligned surrogate model (4-th col) heatmaps demonstrate more diffuse attention spread across the entire object, similar to that of the witness model (ViT-B, 3-rd col), which shows aligned surrogate models learn the common spatial features. For adversarial examples (last four cols), the 5-th and 7-th cols display the heatmaps of adversarial examples generated by the unaligned and aligned surrogate models, respectively. The 6-th and 8-th show the witness model’s responses to these adversarial examples. Notably, the adversarial examples generated by the unaligned surrogate model fail to effectively transfer to the witness model (6-th col) due to still focusing on the target subject, indicating limited cross-model transferability. In contrast, adversarial examples generated by the aligned surrogate model (7-th col) successfully transferred to the witness model (8-th col) as the features are spread out, demonstrating enhanced cross-model transferability achieved through SAA.
@@ 3.1 Experimental Setup
Datasets. Our experiments utilize the ImageNet-compatible dataset ifgsm , a widely adopted subset containing 1,000 images from the ImageNet validation set imagenet . This dataset is commonly used in adversarial robustness studies, such as those in xie2019improving ; mi ; dong2019evading .
Models. To assess the adversarial transferability of different network architectures, we focus on convolutional neural networks (CNNs) and vision transformers (ViTs) as the target models. For CNNs, we select the typically trained ResNet-18 (Res18), ResNet-50 (Res50) and ResNet-101 (Res101) resnet , VGG-19 vgg , DenseNet-121 (DN121) densenet , and Inception-v3 (Inc-v3) inceptionv3 . For ViTs, we evaluate the Vision Transformer (ViT-B) vit , Swin Transformer (Swin-B) swint , Pyramid Vision Transformer (PVT-v2) pvtv2 , and MobileViT-s (MobViT) mobvit .
Metric. Adversarial transferability is quantified by calculating the average attack success rate (Avg. ASR, %) across target models (excluding the surrogate model), with a higher success rate signifying enhanced transferability. In the paper, ’n/a’ is the baseline and defined as the average attack success rate obtained by generating adversarial examples using the surrogate model without any alignment.
Implementation Details. In our experiments, we select the MI mi attack as the baseline for generating adversarial examples with high transferability, as it is widely recognized within the field of adversarial transferability dong2019evading ; xie2019improving ; vmi ; ssa2022 ; wang2021admix ; WANG2024124757 ; sini ; SIA ; sgm ; pna . For MI, we set the perturbation magnitude $\epsilon=16$ mi ; fgsm , perform 10 iterations, with a step size of $\frac{16}{10}=1.6$, and use a momentum $\mu=1$. During the Spatial Adversarial Alignment, all surrogate models are fine-tuned for $1$ epoch using stochastic gradient descent (SGD) with a momentum of 0.9, and no learning rate adjustments are applied. It is important to note that no additional data is used for fine-tuning, as it relies solely on the same training data used for both the surrogate and witness models. The number of adversarial examples generated by SAA is in a $1:1$ ratio with the training dataset. The settings for Model Alignment (MA) ma2025improving are consistent with the parameters specified in the original paper.
@@
Surrogate Witness Attack Target Model
Avg.
ASR (%)
@@
3.2 Performance Comparison
Performance comparison with alignment methods. We first compare with existing alignment methods ma2025improving , where adversarial examples are generated based on MI mi . Table 2 illustrates the performance difference between MA and SAA in terms of adversarial transferability, with SAA demonstrating a significant advantage over MA. For instance, when the surrogate model is Res50, and the witness model is also Res50, SAA achieves a 16.6% improvement in average ASR over the original surrogate model, compared to a modest 3.6% improvement with MA. This highlights that SAA, even without introducing additional information, enhances adversarial transferability through the alignment of adversarial features. Furthermore, when the witness models are DN121, ViT-B, and Swin-B, SAA outperforms MA by 11.9%, 10.4%, and 13.1%, respectively. In addition to the remarkable adversarial transferability that SAA provides, we make two other key observations: (i) MA only considers global features, which makes it difficult to align features between models with large differences, which may lead to a decrease in transferability. When the surrogate model is DN121 and the witness model is Swin-B, the ASR of Vit-B, Swin-B, PVT-v2, and MobViT is not as good as the origin DN121, which shows that relying solely on global features for alignment is not enough, and can only achieve poor transferability. (ii) SAA provides strong cross-architecture transferability. When the surrogate model is Res50, and the witness models are Res50, DN121, ViT-B, and Swin-B, the transferability of SAA on ViTs is improved by 39.06%, 31.29%, 25.51%, and 37.74% respectively compared with MA itself, and it also has high transferability between CNNs.
Table 3: SAA has stronger adversarial transferability with existing transfer attacks.